Circular No. 8676

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2005 P 1 AND S/2005 P 2 M. J. Mutchler, Space Telescope Science Institute; A. J. Steffl, South-

west Research Institute (SwRI); H. A. Weaver, Applied Physics Laboratory, Johns Hopkins University; S. A. Stern, SwRI; M. W. Buie, Lowell Observatory; and W. J. Merline, J. R. Spencer, E. F. Young, and L. A. Young, SwRI, confirm the presence of the two new satellites of Pluto in Hubble Space Telescope ACS/HRC images taken on Feb. 15.6 UT using the same broadband V filter (F606W) employed during the discovery observations in May 2005 (cf. IAUC 8625). Preliminary analyses show that S/2005 P 1 was 2′′.86 from the center of Pluto in p.a. 343^{o} and that S/2005 P 2 was 2′′.03 from Pluto in p.a. 356^{o}; measured V magnitudes are 23.26 ± 0.15 for S/2005 P 1 and 23.7 ± 0.2 for S/2005 P 2. No additional new satellites were detected. For more information, see

http://hubblesite.org/news/2006/09.

V723 CASSIOPEIAE J.-U. Ness and S. Starrfield, Arizona State University; G. Schwarz

and K. Vanlandingham, West Chester University; R. M. Wagner, LBT Observatory; J. Lyke, Keck Observatory; C. E. Woodward, University of Minnesota; D. K. Lynch, The Aerospace Corporation; and J. Krautter, Landessternwarte, Heidelberg-Königstuhl, report that SWIFT XRT observations of V723 Cas (N Cas 1995; IAUC 6213, 6214, 6221) were obtained on Jan. 31.27 UT. V723 Cas was detected in x-rays and exhibited a super-soft-source spectrum with a peak around 0.4 keV. The best derived black-body fit is for a temperature of 340000 K and N_{H} = 1.6 × 10^{21} cm^{−2}. The flux was 7.8 × 10^{−13} erg/cm^{2}/s. This nova is too bright in x-rays to have returned to quiescence, and — given that it has been 11 years since discovery — V723 Cas sets the record for being in outburst longer than any nova observed in x-rays. The previous record holder was GQ Mus, which was observed by ROSAT to be turning off in x-rays nine years after discovery (Shanley et al. 1995, Ap.J. 438, L95). This detection supports the observations of coronal lines in the infrared (IAUC 7259) and suggests that further study at other wavelengths is warranted.

V2575 OPHIUCHI = NOVA OPHIUCHI 2006 N. N. Samus, Russian Academy of Sciences, informs us that N Oph

2006 (cf. IAUC 8671) has been assigned the designation V2575 Oph.

2006 February 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green